OpenStudy (albert0898):

If log3 = x and log5 = y, then log45 is equal to... A) 2x + y B) x^2 + y C) (x^2)(y) D) x + y E) 9x + y The answer is (A). How do you get that?

OpenStudy (anonymous):

How did u get that?

OpenStudy (albert0898):

Sorry I meant (A), not 45 LOL

OpenStudy (anonymous):

LOL, ok well how did come up with that?

OpenStudy (albert0898):

It was in the answers... lol

OpenStudy (anonymous):

Har har har........ Then what do u need help with?

OpenStudy (albert0898):

How was that answer achieved? Like what steps do I need to take?

OpenStudy (tkhunny):

Use the log rules. log(a*b) = log(a) + log(b) log(a^b) = b*log(a) You will also need the Prime Factorization of 45.

OpenStudy (albert0898):

OH! I understand now. Thank you very much @tkhunny

OpenStudy (tkhunny):

Just for the record, for the most part, people in the Chats are not interested in helping anyone do anything. They are there, generally, to play and to make fools of themselves by breaking rules they agreed to follow. This is why they ignored your requests in the chats. There are notable exceptions, but this description certainly describes the norm.

OpenStudy (albert0898):

Yea, I miss the beginning of OpenStudy when people in chat would help. Oh well, it's still just as helpful as it was before. Thanks for the help! :)